# When AIOps Become “AI Oops”:

### v1

@@ 4 Evaluation
We now evaluate the attack methodology described in Section 3. Our aim is to understand its effectiveness across different AIOps environments, and to examine the limits of existing prompt injection defenses in this context. The section is organized as follows. We first describe the experimental setup. We then present the main results, focusing on success rates and the factors that influence them. Finally, we evaluate the performance of established prompt injection defenses under our attack model.
@@ 4.1 Setup
The experiments are built on four axes: the application under test, the agentic framework, the backend LLM, and the adversarial remediation objective. The cross-product of these axes defines the set of scenarios we consider.
@@ 4.1.1 AIOps Agents
Our study uses AIOpsLab [10], an open-source benchmark suite designed to emulate realistic IT operations environments. Compared to earlier benchmarks, AIOpsLab offers a broader set of incident types, agent behaviors, and application architectures.
We consider two agentic frameworks:
  • ReAct [54], which implements tool use with explicit reasoning steps.
  • Flash [58], which augments agent behavior with workflow supervision and the ability to incorporate past failures into future actions. Together these components mimic a self-healing system [26].
@@
Base LLMs
Originally, AIOpsLab implemented the agents relying on GPT-3.5-TURBO and GPT-4-TURBO. We update this configuration with more modern and capable solutions; in particular, we use GPT-4o (gpt-4o-2024-08-06) and GPT-4.1 (gpt-4.1-2025-04-14).
4.1.2 Applications
We evaluate our methodology using two applications implemented in AIOpsLab [10]. Both applications reflect realistic microservices architectures commonly found in production environments.
SocialNet
SocialNet [17] is a social networking platform drawn from the DeathStarBench suite [21]. The system consists of over twenty microservices, including components for user management, content recommendation, media processing, and authentication. Communication between services takes place via HTTP and Thrift APIs, with orchestration provided by Kubernetes. This setup is intended to capture the complexity and heterogeneity of modern cloud-native workloads.
HotelReservation
HotelReservation [1] is a hotel reservation system implemented as a set of loosely coupled services using Go and gRPC. The application maintains both in-memory and persistent state, and includes a recommender module for hotel suggestions. Like SocialNet, all components are deployed on Kubernetes, providing a realistic foundation for the types of operational incidents addressed by AIOps agents.
For the purposes of evaluation, both applications are deployed in a default, fault-free configuration. To simulate incident response, we introduce a metric-based alerting mechanism: an alert is triggered if the system observes more than $N=100$ HTTP errors (specifically, responses with status code 404 or 500) within a 60-second interval. This threshold is chosen to ensure that alerts are generated only in the presence of a sustained anomaly, minimizing spurious activations while maintaining sensitivity to faults. In Appendix B.2, we consider other kinds of possible alerts, such as triggers based on an excessive number of failed logins.
@@ 4.1.3 Malicious Remediations
For each remediation class, we construct an adversarial payload using the adversarial reward-hacking technique described in Section 3.4.1. Payloads are adapted to the context and interfaces of each target application. A complete list of payloads used in the evaluation appears in Table D.1.
@@ 4.2 Attack Results
Figure 8: Attack success over multiple target’s configurations and adversarial objectives. The right-most column reports the number of rounds the agent performed before submitting the solution averaged over the five rounds.
Our evaluation considers all $24$ combinations of application, agent, LLM, and remediation objective. Each configuration is repeated $5$ times, for a total of $120$ independent trials. Before each run, the application environment is reset to its initial state.
We consider an attack successful if the agent produces a remediation that matches the adversary’s intended action. To make this comparison consistent and scalable, we adopt the LLM-as-a-judge methodology [10, 59]: for each trial, we present both the adversary’s intended remediation and the agent’s output to an LLM, using a prompt to determine equivalence (see Appendix D.4). To implement the judge, we use OpenAI’s o4-mini–a model distinct from the one used for the AIOps agent–to reduce potential bias in the evaluation. If the agent does not produce any solution within 35 rounds, we record the trial as a failure.
Results:
@@
Success rates of the attack across each individual setting are reported in Table 8. Figure 9 aggregates the results from Table 8 across the different parameters in the attack setup. Overall, the attack achieves an average success rate of $89.2\%$ across all settings.
Telemetry expressiveness influences exploitability: Among the tested applications, HotelReservation proves substantially more difficult to attack. This is due to the type and format of telemetry the application records. In HotelReservation, the only tainted telemetry an attacker can inject into the system is traces, which are generally less expressive than logs. In this particular case, only the URLs of the requests are stored. Nonetheless, the attack success rate remains meaningfully high, i.e., $82\%$, especially considering that the attacker needs to succeed only once over multiple attempts to compromise the target.
@@
Smarter is safer: Another clear signal from the data is that more advanced agents are more resilient to the attack, particularly when using state-of-the-art base models such as GPT-4.1. The tainted telemetry instances injected by the attacker generally appear structurally broken and include information that has no rational reason to be present (e.g., Figure 7(b)). Advanced models are more likely to detect such inconsistencies and disregard the content, resulting in a failed attack.
Nonetheless, even when using the Flash agent framework–which incorporates reasoning steps and self-adjustment routines–combined with GPT-4.1, the agent still falls for most of the attacks, with an average success rate of 76.7%. Notably, there is no significant variation in attack success rate across different adversarial remediation strategies. This suggests that, despite potential differences in impact, all tested payloads (see Table D.1) are crafted to appear equally plausible, effectively bypassing the agent’s judgment.
@@
Overall, the combination of the telemetry injection technique implemented through AIOpsDoom and the adversarial reward-hacking-base payloads proves to be both effective and reliable in manipulating AIOps agents across diverse settings. In particular, a key factor in the success of the attack is the use of adversarial reward-hacking. This is evidenced in Appendix A, where baseline prompt injection attacks, under identical target and telemetry injection conditions, yield a 0% attack success rate, in contrast to the 89.2% success rate achieved via adversarial reward-hacking.
Additional results obtained under different configurations are presented in Appendix B.1.
4.3 Evaluation of Prompt Injection Defenses
Figure 10: Evasion rate of the proposed attack against three prompt injection defenses.
A central question in this work is whether existing prompt injection defenses are effective against the class of attacks described above. In particular, we are interested in whether deliberate reward-hacking payloads can evade systems designed to detect adversarial prompts in large language model inputs.
Results:
Figure 10 reports the average evasion rate of the adversarial reward-hacking payloads across all the runs for the three detectors. In the plot, evasion rate of 100% means that no adversarial reward-hacking payloads have been classified as malicious by the detector. In our evaluation, adversarial reward-hacking payloads achieved perfect evasion success rate against two of the three detectors: Prompt-Guard2 and PromptShields. The most performant defense turns out to be DataSentinel, where the attack has 85% evasion rate; that is, 15% recall. Despite outperforming the others, DataSentinel still failed to provide adequate protection against the attack.
Why does it work?
@@ 5 Securing AIOps
In this section, we propose a novel defense mechanism called AIOpsShield, which leverages the unique features of the AIOps setting to nullify all adversarial inputs.
@@ 5.1 AIOpsShield
Despite many proposals by the academic community [14, 48, 31, 9, 4], there is still no (reliable) solution for prompt injection that works consistently in all contexts. In the general case, especially against adaptive adversaries [5], it continues to be an open problem. As demonstrated in our experiments (Section 4.3), even state-of-the-art and industry-level detectors often fail to block attacks that deviate from expected patterns. Moreover, many existing prompt injection defenses are not easily generalizable to open-ended, multi-stage, and loosely defined tasks, such as incident response and root cause analysis.
@@
In this work, we demonstrate that this holds true for the AIOps setting. Notably, this environment presents a set of inherent properties that enable defenders to address adversarial inputs in a simple and effective manner. The key properties of AIOps’s tasks that allow this are:
@@
    User-provided data offer only limited utility in solving incident response (a claim that we verify in Section 5.2) and can be safely excluded without impacting the agent’s overall functionality. [^5] In contrast, untrusted data–such as web content–plays a critical and irreplaceable role in general LLM applications. Removing such data would fundamentally alter the application’s functionality.
@@ 5.1.1 AIOpsShield: Setup Phase
Once the endpoints have been collected, AIOpsShield runs the AIOpsDoom’s fuzzer and sets the payload to a unique string (hereafter, “CANARY”) which serves as a canary for taint analysis [35]. After fuzzing is complete, AIOpsShield extracts all logs, metrics, and traces from the application via querying the observability stack across all available namespaces and scopes. It then parses the output to identify any telemetry instances that contain the canary string (including small variations, such as case-insensitive matches or base64-encoded versions). An example tainted log entry is shown in Figure 11(a). By property $\spadesuit$, these entries capture all attacker-controlled vectors for injecting payloads into the observability stack. Their completeness depends on the coverage of the preceding endpoint enumeration.
@@
Figure 11: Example of a tainted telemetry abstraction and template derivation process for AIOpsShield. (a) Example of an error log generated by the SocialNet application in [10] (an unfollow request submitted for a non-existent user), triggered by the fuzzer. (b) Regex automatically generated to match and parse the log in (a). (c) List of parameters extracted by the regex, along with the assigned contextual labels.
@@
To automate template generation, we use an LLM-based approach. Given a tainted telemetry entry (e.g., Figure 11(a)), the LLM: (1) generates a robust regular expression to match the entry and extract variable parameters (e.g., timestamps or user inputs); an example is shown in Figure 11(b), (2) assigns semantically meaningful labels to each parameter (Figure 11(c)) to aid abstraction. Parameters containing the canary string are marked as untrusted. In Figure 11(b), for instance, only the “username” field is considered tainted. All other parameters are treated as trusted and are not abstracted at inference time. The regex extraction process is designed to be robust against evasion. Implementation details are in Appendix C.
If the telemetry entry is strictly structured (e.g., JSON), as is common with metrics and traces, step (1) can be skipped. Instead, we algorithmically derive a schema from the tainted telemetry and use it to match similar structured entries, eliminating the need for regular expressions. Only step (2) (parameter labeling) is performed in this case.
Remarks. Note that telemetry data used the to derive the templates at this stage do not contain any external users’ inputs; thus, the result of the setup phase cannot be manipulated by attackers within any realistic threat model. [^6] If an attacker can compromise this step, it implies they already have control over the application before the application even starts.
@@ 5.1.2 AIOpsShield: Runtime Phase
Figure 12: Example of application of AIOpsShield at inference time. The first panel shows raw, tainted telemetry from the application. The second panel displays the resulting sanitized output.
@@
Abstracting and Mapping Untrusted Inputs. AIOpsShield maintains a consistent variable scope for untrusted parameters by assigning each instance encountered at inference time a unique abstract name. If a parameter has been seen before, it is linked to its existing name; otherwise, a new name is generated by combining the setup-time label with a unique identifier. For example, “Mozilla/5.0 (Macintosh;...” might be mapped to USERAGENT-42, where 42 is a counter or randomly generated ID. This variable scope is global and consistent across runs, ensuring that parameters (whether malicious or honest) always map to the same abstract name, both within a single run and across multiple runs.
Once the untrusted parameters have been replaced with their abstract representations, the telemetry entry is reassembled according to the template. This generates a sanitized telemetry entry that is returned to the agent in place of the original one. This operation is fully transparent to the agent, which continues its operation unaffected.
On Dereferencing Abstract Variables. Once the agent run is complete, any abstract variable names present in the produced root cause analysis/remediation output from the agent (if any) are mapped back to their original values. However, the defender must ensure that these outputs are not then used as input to another agent (e.g., an automated remediation agent) to avoid second-order attacks. If this is the case, dereferencing should be disabled.
@@ 5.1.3 On the Effectiveness of AIOpsShield
To test the implementation of AIOpsShield, we run the automated setup phase of the tool (see Section 5.1.1) on the applications SocialNet and HotelReservation, resulting in 84 and 12 templates, respectively. Then, we rerun the attacks of Section 4.2 but apply AIOpsShield on the agents. None of the attacks result in success; every injected payload is sanitized by AIOpsShield before reaching the agent.
@@
Limitations and Generalizability of AIOpsShield. AIOpsShield fully prevents the attacks introduced in Section 3 as well as any form of direct adversarial input via telemetry, regardless of the nature of the adversarial input (assuming a sound implementation of AIOpsShield and setup). However, it is not possible to exclude that telemetry injection is the only attack vector adversaries can exploit in order to manipulate AIOps agents.
@@
5.2 AIOpsShield’s Impact on Utility
AIOpsShield manipulates telemetry data by abstracting untrusted inputs. This unavoidably reduces the information that a sanitized telemetry entry carries, potentially limiting the agent’s ability to solve the underlying AIOps task. In this section, we empirically show that this is not the case in practice. To do so, we run the agents implementing AIOpsShield on the AIOps benchmark AIOpsLab [10], and show that their performance remains unaltered compared to the same agents not relying on AIOpsShield.
Setup. In this setting, we focus on testing the Flash agent, as it has been shown to be the most effective at solving the tasks in the original work [10]. In our setup, we use GPT-4.1 as the base LLM for the agent. For the evaluation, we consider 12 different fault scenarios listed in Table 1. We then run the Flash agent with and without AIOpsShield and measure its average success rate across the 12 scenarios. We repeat each run 3 times. Note that no attack is carried out here. The objective of these evaluations is to verify that the agent preserves utility when working on telemetry sanitized via AIOpsShield.
Results. The success rates of the agents with and without AIOpsShield are shown in Table 1. Both agents perform nearly identically across all tasks, with an average success rate of around 50%. The only difference is that the agent with AIOpsShield fails one additional run (last row) compared to the agent without AIOpsShield. Even if this failure is attributed to the use of AIOpsShield rather than randomness, the utility loss remains negligible.
AIOpsLab [10] tasks: w/o w/
user_unregistered_mongodb-analysis-2 0 / 3 0 / 3
k8s_target_port-misconfig-mitigation-3 2 / 3 2 / 3
k8s_target_port-misconfig-detection-3 3 / 3 3 / 3
k8s_target_port-misconfig-mitigation-2 0 / 3 0 / 3
k8s_target_port-misconfig-analysis-2 0 / 3 0 / 3
user_unregistered_mongodb-localization-2 0 / 3 0 / 3
user_unregistered_mongodb-detection-2 3 / 3 3 / 3
k8s_target_port-misconfig-localization-3 3 / 3 3 / 3
k8s_target_port-misconfig-analysis-3 0 / 3 0 / 3
user_unregistered_mongodb-mitigation-2 2 / 3 2 / 3
k8s_target_port-misconfig-localization-2 3 / 3 3 / 3
k8s_target_port-misconfig-detection-2 2 / 3 1 / 3
Table 1: Results of a Flash agent based on GPT-4.1 on the 12 tasks from the AIOpsLab benchmark without (w/o) and with (w/) AIOpsShield. Each task is repeated 3 times for the agent and the number of successful runs is reported in the table.

### v2

@@ 4 Evaluation
We now systematically evaluate the attack methodology described in Section 3. Our aim is to understand its effectiveness across different AIOps environments and applications. The section is organized as follows. We first describe the experimental setup. We then present the main results, focusing on success rates and the factors that influence them.
@@ 4.1 Setup
The experiments are built on four axes: the agentic framework used to implement the AIOps agent, the backend LLM used to run the agent, the system/application the AIOps agent is deployed on ($\hat{t}$), and the adversarial remediation objective chosen by the attacker. The cross-product of these axes defines the set of scenarios we consider
@@ 4.1.1 AIOps Agents
Our study uses AIOpsLab [10]; the leading benchmark suite designed to emulate realistic IT operations environments. AIOpsLab offers a broader set of incident types, agent behaviors, and application architectures. We experiment with the two most performant [10] agents provided in AIOpsLab:
  • ReAct: A ReAct-based agent [54].
  • Flash: A Flash-based agent [58], which augments tool usage with workflow supervision and the ability to incorporate past failures into future actions.
@@
Base LLMs Originally, AIOpsLab implemented those agents relying on GPT-3.5-TURBO and GPT-4-TURBO. We update this configuration with more modern and capable solutions; in particular, we use GPT-4o (gpt-4o-2024-08-06) and GPT-4.1 (gpt-4.1-2025-04-14).
4.1.2 System / Applications
To model the infrastructure on which the AIOps agent is deployed ($\hat{t}$), we use all three applications provided in AIOpsLab [10]. Each of these applications reflects realistic microservice architectures commonly found in production environments. Those are:
SocialNet: SocialNet [18] is a social networking platform drawn from the DeathStarBench suite [22]. The system consists of over twenty microservices, including components for user management, content recommendation, media processing, and authentication. Communication between services takes place via HTTP and Thrift APIs, with orchestration provided by Kubernetes.
AstronomyShop: AstronomyShop [12] implements an e-commerce platform with functionality for cart management, product catalog browsing, checkout, and payment processing. The system comprises over a dozen services written in languages such as Go, Python, Java, and JavaScript, communicating via HTTP and gRPC.
HotelReservation: HotelReservation [1] is a hotel reservation system implemented as a set of loosely coupled services using Go and gRPC. The application maintains both in-memory and persistent state, and includes a recommender module for hotel suggestions.
For the purposes of evaluation, applications are deployed in a default, fault-free configuration. To simulate incident response, we introduce a metric-based alerting mechanism: an alert is triggered if the system observes more than $N=100$ HTTP errors (specifically, responses with status code 404 or 500) within a 60-second interval. This threshold is chosen to capture a realistic setting where alerts are generated in the presence of a sustained anomaly. Nevertheless, this value can be arbitrarily set, as it does not directly affect the attack success (as long as it is set to a realistic value). In Appendix C, we consider other kinds of possible alerts, such as triggers based on an excessive number of failed logins as well as no alert at all.
@@ 4.1.3 Malicious Remediations
We selected these remediations to represent a broad and diverse set of realistic attack objectives, spanning different levels of severity. For each remediation class, we automatically construct an adversarial reward-hacking payload using the pipeline described in Appendix B. Therefore, the payloads are adapted to the context and interfaces of each target application. The complete list of payloads used in the evaluation appears in Table G.1 in Appendix G.
@@ 4.2 Attack Results
Our evaluation considers all $36$ combinations of application, agent, LLM, and adversarial remediation objective. Each configuration is repeated $5$ times, for a total of $180$ independent trials. Before each run, the application environment is reset to its initial state and re-deployed on fresh VMs.
We consider an attack successful if the agent produces a remediation that matches the adversary’s intended action. To make this comparison consistent and scalable, we adopt the LLM-as-a-judge methodology [10, 59]: for each trial, we present both the adversary’s intended remediation and the agent’s output to an LLM, using a prompt to determine equivalence (see Appendix G.5). To implement the judge, we use OpenAI’s o4-mini–a model distinct from the one used for the AIOps agent–to reduce potential bias in the evaluation. If the agent does not produce any solution within 35 rounds, we record the attack as a failure.
Results: Figure 8 reports aggregated attack success rate across the different setup parameters.
@@
Table 1: Attack success over multiple target’s configurations and adversarial objectives. The right-most column reports the number of rounds the agent performed before submitting the solution averaged over the five rounds.
@@
We provide fine-grained attack outcomes across each individual setting in Table 1. Overall, the attack achieves an average success rate of $90\%$ across all settings.
Telemetry expressiveness influences exploitability: Among the considered axes, the most prominent factor impacting the attack success rate is the application, with HotelReservation being the hardest to attack. To investigate the cause, we perform a fully white-box, post-mortem analysis of the three applications following the attacks.
The lower success rate on HotelReservation is due to the type and format of inducible tainted telemetry. HotelReservation represents an edge case: the only telemetry an attacker can inject into the system consists of traces, recording only the URLs of the requests. As a result, the payload is heavily distorted by URL encoding. Nonetheless, the attack success rate remains meaningfully high, i.e., $82\%$, especially considering that the attacker needs to succeed only once over multiple attempts to compromise the target. Other applications offer increasingly diverse telemetry injection points compared to HotelReservation, giving attackers greater flexibility in delivering their payloads. The lower performance of AstronomyShop relative to SocialNet is primarily due to the fact that AstronomyShop monitors a consistently larger pool of non-taintable telemetry, which reduces the likelihood that the agent will actually access the payload during execution.
@@
Smarter” is safer: Another signal from the data is that more advanced agents are more resilient to the attack, particularly when using state-of-the-art base models such as GPT-4.1. The tainted telemetry instances injected by the attacker generally appear structurally broken and include information that has no rational reason to be present (e.g., Figure 7(b)). Advanced models such as GPT-4.1 are more likely to detect such inconsistencies and disregard the content, resulting in a failed attack. Nonetheless, even when using the Flash agent framework (which incorporates reasoning steps and self-adjustment routines) combined with GPT-4.1, the agent still falls for most of the attacks, with an average success rate of 82.2% (see Table 1).
[…]
@@
Overall, the combination of the telemetry injection technique implemented through AIOpsDoom and the adversarial reward-hacking-base payloads proves to be reliable in manipulating AIOps agents across diverse settings. In particular, a key factor in the success of the attack is the use of adversarial reward-hacking. This is evidenced in Appendix A, where baseline prompt injection attacks, under identical target and telemetry injection conditions, yield a 0% attack success rate, in contrast to the 90% success rate achieved via adversarial reward-hacking. Additional results obtained under different configurations are presented in Appendix C. In Appendix C.1, we test the attacks against a no-alert deployment (see Section 3.4.2), where the agent operates on a fixed schedule rather than being activated by a specific alert/ticket. The results are consistent with the observations reported above.
Adversarial Reward-Hacking bypasses prompt injection detectors: Another interesting property of adversarial reward-hacking is that, due to its semantic differences from standard prompt injection attacks, it is able to consistently evade state-of-the-art defenses such as PromptShields [33], Prompt-Guard2 [6], and DataSentinel [31]. A detailed empirical analysis and further discussion of this phenomenon are provided in Appendix D.
@@ 5 Securing AIOps
Introduced the attack, in this section, we propose a novel defense mechanism called AIOpsShield, which leverages the unique features of the AIOps setting to nullify injection via telemetry data.
@@ 5.1 AIOpsShield
Despite many proposals by the academic community [15, 48, 31, 9, 4], there is still no (reliable) solution for prompt injection that works consistently in all contexts. In the general case, especially against adaptive adversaries [5], it continues to be an open problem. As demonstrated in our experiments (Section D), even state-of-the-art and industry-level detectors often fail to block attacks that deviate from expected patterns. Moreover, most existing prompt injection defenses that rely on policy enforcement or expected task behavior [15, 4] are not easily generalizable to open-ended, multi-stage, and loosely defined tasks such as incident response and root cause analysis.
@@
Next, we demonstrate that this holds true for the AIOps setting. Notably, this environment presents a set of inherent properties that enable defenders to address adversarial inputs in a simple and effective manner. The key properties of AIOps’s tasks that allow this are:
@@
    User-provided data offer only limited utility in solving incident response and RCA (a claim that we verify in Section F) and can be safely excluded without impacting the agent’s overall functionality. [^6] In contrast, untrusted data–such as web content–plays a critical and irreplaceable role in general LLM applications e.g., Q&A on a provided corpus. Removing such data would fundamentally alter the application’s functionality.
@@ 5.1.1 AIOpsShield: Setup Phase
Once the endpoints have been collected, AIOpsShield runs the AIOpsDoom’s fuzzer and sets the payload to a unique string (hereafter, “CANARY”) which serves as a canary for taint analysis [35]. After fuzzing is complete, AIOpsShield extracts all logs, metrics, and traces from the application via querying the observability stack across all available namespaces and scopes. It then parses the output to identify any telemetry instances that contain the canary string (including small variations, such as case-insensitive matches or base64-encoded versions). An example tainted log entry is shown in Figure 9(a). By property $\spadesuit$, these entries capture all attacker-controlled vectors for injecting payloads into the observability stack. Their completeness depends on the coverage of the preceding endpoint enumeration and fuzzing.
@@
Figure 9: Example of tainted telemetry abstraction and template derivation for AIOpsShield. (a) Error log from SocialNet in [10] (unfollow request for a non-existent user), triggered by the fuzzer. (b) Regex generated to match and parse the log in (a). (c) Extracted parameters with contextual labels.
@@
To automate template generation, we use an LLM-based approach. Given a tainted telemetry entry (e.g., Figure 9(a)), the LLM: (1) generates a robust regular expression to match the entry and extract variable parameters (e.g., timestamps or user inputs); an example is shown in Figure 9(b), (2) assigns semantically meaningful labels to each parameter (Figure 9(c)) to aid abstraction. Parameters containing the canary string are marked as untrusted. In Figure 9(b), for instance, only the “username” field is considered tainted. All other parameters are treated as trusted and are not abstracted at inference time. The regex extraction process is designed to be robust against evasion. Implementation details are in Appendix E. If the telemetry entry is strictly structured (e.g., JSON), as is common with metrics and traces, step (1) can be skipped. Instead, we algorithmically derive a schema from the tainted telemetry and use it to match similar structured entries, eliminating the need for regular expressions. Only step (2) (parameter labeling) is performed in this case.
Remarks. Note that telemetry data used to derive the templates at this stage do not contain any external users’ inputs; thus, the result of the setup phase cannot be manipulated by attackers within any realistic threat model. [^7] If an attacker can compromise this step, it implies they already have control over the application before the application even starts.
@@ 5.1.2 AIOpsShield: Runtime Phase
Figure 10: Example of application of AIOpsShield at inference time. The first panel shows raw, tainted telemetry. The second panel displays the resulting sanitized output.
@@
Abstracting and Mapping Untrusted Inputs.
AIOpsShield maintains a consistent variable scope for untrusted parameters by assigning each instance encountered at inference time a unique abstract name. If a parameter has been seen before, it is linked to its existing name; otherwise, a new name is generated by combining the setup-time label with a unique identifier. For example, “Mozilla/5.0...” might be mapped to USERAGENT-42, where 42 is a counter or randomly generated ID. This variable scope is global and consistent across runs, ensuring that parameters (whether malicious or honest) always map to the same abstract name, both within a single run and across multiple runs.
Once the untrusted parameters have been replaced with their abstract representations, the telemetry entry is reassembled according to the template: Figure 10 (bottom). This generates a sanitized telemetry entry that is returned to the agent in place of the original one. This operation is fully transparent to the agent, which continues its operation unaffected. Furthermore, because sanitization occurs on the fly on the agent side, the raw telemetry retains all available information, thereby still enabling fine-grained inspection by human operators.
@@ 5.1.3 On the Effectiveness of AIOpsShield
To test the implementation of AIOpsShield, we run the automated setup phase of the tool (see Section 5.1.1) on the three applications SocialNet, HotelReservation, and AstronomyShop resulting in 84, 12, and 132 templates, respectively. Then, we rerun the attacks of Section 4.2 but apply AIOpsShield on the agents. None of the attacks result in success; every injected payload is sanitized by AIOpsShield before reaching the agent.
@@
Limitations and Generalizability of AIOpsShield. AIOpsShield prevents the attacks introduced in Section 3. However, it is not possible to exclude that telemetry injection is the only attack vector adversaries can exploit in order to manipulate AIOps agents.
@@
Impact on Utility: AIOpsShield manipulates telemetry data by abstracting untrusted inputs. This unavoidably reduces the information that a sanitized telemetry entry carries, potentially limiting the agent’s ability to solve the underlying AIOps task. In Appendix F, we empirically show that this is not the case in practice.